The minimum force required to start pushing ab body up a rough inclined plne is F1 while the minimum foce needed to prevent it from sliding down is F2. If the inclined plane makes an angle A with the horizontal such that tanA=2mu,the the ratio F1/F2 is ??

Grade:11

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
If we write the forces, we get:
F_{1}=mg\left ( sin\theta+\mu cos\theta \right ); F_{2}=mg\left ( sin\theta-\mu cos\theta \right )
Hence, using this we get:
2sin\theta=\frac{1}{mg}\left ( F_{1}+F_{2} \right );2\mu cos\theta=\frac{1}{mg}\left ( F_{1}-F_{2} \right )
thus, we get:
F_{1}=3F_{2}
Regards
Sumit